OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 9 * 72 = 9* _____ + 9 * 2 = ______

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

idk what to do, i did the problem 4 times already

OpenStudy (igreen):

Is this the problem? \(9 \times 72 = 9 \times~? + 9 \times 2\)

OpenStudy (anonymous):

yes, but at the end of to, somthing has to equel somthing

OpenStudy (anonymous):

at the end of 2

OpenStudy (igreen):

Oh, okay..so it's splitting \(9 \times 72\) into two problems..and it gave you \(9 \times 2\), so to find the other number we have to multiply we can just subtract 72 - 70.

OpenStudy (anonymous):

we come out with 2

OpenStudy (igreen):

Oh, sorry, I meant 72 - 2

OpenStudy (anonymous):

oh then we come out with 70

OpenStudy (anonymous):

so the missing number is 70?

OpenStudy (igreen):

Yes, so we have: \(9 \times 72 = 9 \times 70 + 9 \times 2 =~?\) Just multiply 9 * 72 to get the last number.

OpenStudy (anonymous):

648, omg thanks

OpenStudy (anonymous):

so were done? right?

OpenStudy (igreen):

Yep! Our final answer is: \(9 \times 72 = 9 \times 70 + 9 \times 2 = 648\)